OpenStudy (anonymous):

Graph the inequality. 4x + 6y 10

OpenStudy (kymber):



hero (hero):

In order for someone to help you with this, you'll need to indicate: \[<,>, \le \text{or}\ge\]

OpenStudy (anonymous):

\[\ge\]

hero (hero):

1. Original Problem \[4x + 6y \ge 10\] 2. Isolate y on the left side of the inequality: \[y \ge \frac{10-4x}{6}\] 3. Graph the inequality: